OpenStudy (anonymous):

the length of line segment CB is 64 units and the length of line segment BG is 137 units. What is the length of line segment EA?

OpenStudy (anonymous):

Please draw the required figure.

OpenStudy (anonymous):

any ideas?

OpenStudy (anonymous):

Its problem based on trigonometry. Familiar with that?

OpenStudy (anonymous):

64 sqrt 2 73sqrt 3 128 146

OpenStudy (anonymous):

those are the answers

OpenStudy (anonymous):

Yup.

OpenStudy (anonymous):

any idea on which one?

OpenStudy (anonymous):

oh. I thought u worked it out. Anyways, First find BA BA = CB / tan45 BA = CB = 64 AG = BG - BA AG = 137- 64 = 73 cos(60) = AG / AE AE = AG / cos 60 AE = 2 * AG = 2*73 therfore AE = 146

OpenStudy (anonymous):

what about this one? A toddler’s playground has two slides on either side. The slide on the left is situated at a 45° angle to the ground and is 64 centimeters away from the base of the playground. The slide on the right is positioned at a 30° angle to the ground and is 106 centimeters away from the playground’s base. What is the total length of both slides on this playground rounded to the nearest tenth of a centimeter?

OpenStudy (anonymous):

any ideas?

OpenStudy (anonymous):

tan 45 = 1 right! So, the length of center support should be equal to 64. Now use Pythagorus theorem to find out length of the slides.

OpenStudy (anonymous):

what is the total length of the slides?

OpenStudy (anonymous):

Add the two values u get after using pythagorus theorem. Are u familiar with the theorem?

OpenStudy (anonymous):

what is the yellow slides hypotenuse?